Operations Report – January 21st

Crew 203 Operations Report 21-JAN-2019

SOL: 9

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 5hr 00min; Turned on last night (20Jan2019) at 17:40; Turned off about 20:40; It was shuted down all night long; This morning (21Jan2019) it was Turned On from 9:00 to 11:00, it was running nominally. At 14:00 the SOC was almost 60%, it was a pretty cloudy day with some snow shower, so the Generator was turned On. It ran for 2 hours and the SOC display was 100%. It wasTurned On again tonight (21Jan2019) at 17:40 and is running nominally.

Solar SOC – Turned on (20Jan2019) 85%; Turned off (21Jan2019) 50%; Turned on (21Jan2019) 86%

Diesel Reading – Still going under 1/2 tank.

Propane Reading – Almost 20%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 40%; 215 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – Over 49%; 162 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:30 to refill tank

Water Meter: 01403355

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Used for EVA to the Moons.

Hours: 69.6

Beginning charge: 100%

Ending charge: 38%

Currently charging: Yes

Opportunity rover used: Used to test brakes.

Hours: 46.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Used for EVA to the Moons.

Hours: 71.5

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 still pending for review. Suit #11 quit during EVA. Currently charging.

Summary of Hab operations: Today we realized that half of the toilet paper provided is actually kitchen towels. 🙁

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.